# HUME: Measuring the Human-Model Performance Gap in Text Embedding Tasks

### v2

@@ 4.3 Dataset Quality and Evaluation Challenges
Our analysis reveals systematic quality issues in several MTEB datasets that fundamentally compromise their reliability as evaluation benchmarks. Human performance variation often correlates with underlying ambiguity rather than genuine human limitations, providing a diagnostic tool for identifying problematic evaluation targets. Below we highlight two such examples:
@@
These patterns suggest that apparent “superhuman” model performance often occurs precisely where human agreement is lowest, indicating that models excel not through superior understanding but through consistent reproduction of systematic labeling patterns. This raises concerns about the label quality in embeddings benchmarks, and we encourage future benchmark developers to critically examine the datasets before including them in a benchmark, potentially using human annotations framework like HUME. Detailed analysis of specific quality issues is provided in Appendix C.
@@ 5.1 Task Quality and Embedding Evaluation Reliability
Models achieve their highest relative performance precisely where human experts show the least agreement. This striking pattern raises fundamental questions about what constitutes reliable evaluation in embedding benchmarks.
Tasks with high human agreement, such as reranking ($\rho=0.64$$0.85$) and toxicity classification ($\kappa=0.55$), provide reliable evaluation targets, while low-agreement tasks like emotion classification ($\kappa=0.39$) and ArXiv clustering (ARI = 0.001-0.001) suffer from ambiguous guidelines or subjective judgments. While some cases involve clear quality issues, the mechanisms behind this inverse relationship remain unclear. Models may excel through superior semantic understanding or by exploiting systematic annotation patterns.
Cultural factors further complicate evaluation. Humans retain substantial advantages in Arabic semantic similarity (67.5% vs. 40.9% best model) and multilingual sentiment analysis, revealing genuine model limitations in cross-cultural understanding.
These findings suggest reliable evaluation depends as much on task quality as model capability. Rather than automatically treating high model performance as progress, we recommend prioritizing high-agreement tasks for model and benchmark development, addressing cultural competence gaps, and critically examining whether apparent model superiority on ambiguous tasks reflects genuine capabilities or systematic biases in evaluation datasets.
@@ 5.2 Implications for Model Development and Evaluation Practices
Address Cultural and Linguistic Competence Gaps: The substantial human advantages in non-English tasks reveal critical model limitations that scaling training data alone cannot address. Arabic semantic similarity shows the largest human advantage (67.5% vs. 40.9% best model), while multilingual sentiment analysis demonstrates consistent human superiority in non-English languages (95.0% Arabic, 92.5% Russian). These gaps suggest that current models lack the cultural and contextual knowledge necessary for cross-lingual understanding, requiring architectural innovations or training approaches that go beyond simple data scaling to capture cultural nuances and contextual understanding.
@@
Replace Problematic Benchmark Datasets: Our analysis identifies specific datasets that compromise benchmark reliability and should be replaced in future MTEB iterations: emotion classification ($\kappa=0.39$), ArXiv clustering (ARI=0.001ARI=-0.001), and STS22-Russian (systematic parsing artifacts). These tasks provide unreliable evaluation targets that may mislead model development efforts by rewarding pattern matching to flawed gold standards. Replacement datasets should demonstrate reasonable human agreement and clear task specifications, validated through human evaluation studies before inclusion in benchmark suites.
Develop Culturally-Aware Evaluation Frameworks: Current benchmarks may underestimate cross-cultural understanding challenges by focusing on tasks where models can leverage English-centric training advantages. The pronounced human advantages in Arabic and other non-English tasks demonstrate the need for evaluation frameworks that properly assess cultural and contextual understanding. Future benchmarks should include more diverse cultural contexts and ensure that high model performance reflects genuine cross-cultural competence rather than exploitation of training data biases.
Consider Agreement-Weighted Evaluation: Model performance should be interpreted in light of human agreement levels, with evaluation practices adapted accordingly. We propose reporting model performance alongside human agreement metrics to provide proper interpretive context. A model achieving 85% accuracy on emotion classification (185% of human performance, $\kappa=0.39$) represents a fundamentally different achievement than 85% on reranking (97% of human performance, $\rho=0.75$). High model performance on low-agreement tasks should be viewed skeptically as potential artifacts of flawed evaluation targets rather than genuine capability improvements.
5.3 Limitations and Future Directions
Our study has several limitations. First, while three annotators participated overall, only two annotations were collected for most tasks, constraining our ability to fully characterize human judgment distributions. Sample sizes (20–50 instances per task) balance feasibility and reliability but may not capture the full complexity of human performance variation. Additionally, annotators were “average” or above-average raters without specialized training; expert annotators could perform better, particularly on technical tasks like academic clustering. However, we believe that the samples provided clearly show that decisions on evaluation datasets can be made from even a few samples.
Future work should expand annotator pools to better capture human judgment variability and extend evaluations to additional MTEB task categories. Crucially, the field needs rigorous research on designing tasks that balance clear specification with meaningful challenge, avoiding both ambiguity, which depresses human agreement, and oversimplification, which diminishes the evaluation’s ability to discriminate model capabilities.
Finally, embedding evaluation should aim not for model “superhuman” performance but for frameworks that reliably distinguish genuine semantic understanding from sophisticated pattern matching. Achieving this requires careful attention to task quality, human agreement, and the cultural contexts shaping language understanding.
@@ 6 Conclusion
We introduce HUME, a comprehensive human evaluation framework for MTEB, addressing a critical gap in understanding empirical performance bounds for embedding models. By measuring human performance across 16 datasets spanning reranking, classification, clustering, and STS, we establish baselines that reframe how model achievements should be interpreted.
Our findings show that human performance varies substantially by task categories. Tasks with high agreement provide reliable benchmarks, while low-agreement tasks often indicate design issues rather than human limitations. Apparent model “superhuman” performance tends to coincide with low human agreement, suggesting models exploit patterns rather than achieving true semantic understanding.
HUME provides both a methodology and baseline measurements for evaluating embedding models. By releasing our resources and protocols, we aim to support the community in adopting human-grounded evaluation practices and developing models better aligned with human judgment and real-world language understanding.
Appendix A Detailed Results by Task Category
@@
Appendix B Task Examples
@@
Appendix C Task Quality Analysis
C.1 Dataset Quality Issues
@@ C.1.1 STS22-Russian
C.1.2 Multilingual sentiment classification-Russian
@@
  • Similar to the issues described in § C.1.1, the dataset contains parsing artifacts from web pages.
C.1.3 Emotion Classification
@@
C.1.4 ArXiv Clustering Challenges
@@ C.2 Impact on Evaluation
Appendix D Inter-Annotator Agreement Analysis
D.1 Agreement Metrics by Task Category
@@ D.1.4 STS Tasks
D.2 Agreement Patterns and Task Reliability
D.2.1 High-Agreement Tasks (Reliable Benchmarks)
@@
D.2.2 Low-Agreement Tasks (Problematic Benchmarks)

### v3

@@ 4.2 Cross-Lingual Performance Analysis
Human win rates vary substantially by language and comparison baseline (see Appendix G for detailed breakdown). Against the best models, humans win only 15% of tasks overall, but this rises to 62% against median models. The advantage is strongest for non-English tasks: humans achieve a 29% win rate on multilingual tasks versus 0% on English-only tasks against best models.
@@ 4.3 Dataset Quality and Evaluation Challenges
Our analysis reveals systematic quality issues in several MTEB datasets that fundamentally compromise their reliability as evaluation benchmarks. Human performance variation often correlates with underlying ambiguity rather than genuine human limitations, providing a diagnostic tool for identifying problematic evaluation targets. See Appendix D for further qualitative analysis of these failure modes; below, we highlight two such examples.:
@@
These patterns suggest that apparent “superhuman” model performance often occurs precisely where human agreement is lowest, indicating that models excel not through superior understanding but through consistent reproduction of systematic labeling patterns. This raises concerns about the label quality in embeddings benchmarks, and we encourage future benchmark developers to critically examine the datasets before including them in a benchmark, potentially using human annotations framework like HUME. Detailed analysis of specific quality issues is provided in Appendix D.
4.4 Can LLMs Replace Human Annotators?
We benchmark nine LLMs as annotators to assess whether they can serve as reliable proxies for human judgment. As shown in Table 2, the best-performing LLM (GPT-4.1-mini) achieves 76.1% average accuracy, falling short of human performance (81.2%). We exclude clustering tasks in this comparison due to the difficulty of eliciting cluster assignments from generative models.
Task-specific patterns reveal important limitations. On classification, LLMs approach human performance (GPT-5: 78.9% vs. Human: 79.1%). However, a substantial gap emerges in reranking, where humans achieve 88.3% compared to the best LLM at 78.0% (Mistral-Small): a 10-point deficit on tasks where humans show high agreement ($\rho=0.64$$0.85$). For STS, humans (76.5%) outperform all LLMs (best: 75.0%, Mistral-Small). Embedding models achieve the highest scores across all categories (Classification: 80.3%, Clustering: 79.1%, Reranking: 94.8%, STS: 77.1%). Detailed per-task LLM performance is provided in Appendix I.
To assess whether humans and LLMs face similar challenges, we computed Spearman rank correlations between human and LLM performance across the 19 task-language pairs. The moderate positive correlation ($\rho=0.52$, $p<0.05$, $n=19$) indicates that tasks where humans perform well also tend to be tasks where LLMs perform well, suggesting partially shared difficulty patterns. However, the correlation is moderate rather than strong, indicating that humans and LLMs do not face identical challenges across all tasks (see § I.1 for detailed correlation analysis).
Task Human GPT-5 GPT-4.1 Gemini Mistral Qwen3 Best Emb.
Full Mini Full Mini 2.5 Flash Small-24B-I 30B 32B Coder Score
Classification 79.1 78.9 77.2 76.6 76.1 77.6 73.8 74.2 73.0 76.3 80.3 (jasper)
Reranking 88.3 75.1 75.5 75.7 77.2 76.2 78.0 75.6 74.8 73.8 94.8 (e5)
STS 76.5 73.0 69.0 74.9 74.9 69.3 75.0 67.1 68.6 71.3 77.1 (jasper)
Average 81.2 75.8 74.1 75.8 76.1 74.5 75.5 72.4 72.2 73.9
Table 2: LLM-as-annotator performance compared to human annotators and best embedding models per task category. Human and LLM performance is computed over 19 task-language pairs (clustering tasks excluded due to difficulty eliciting cluster assignments). Best embedding model per task category shown with abbreviated name: jasper (jasper_en_vision_language_v1), SFR (SFR-Embedding-Mistral), e5 (multilingual-e5-large). Bold indicates best LLM performance (humans and embedding models consistently outperform LLMs and are not bolded).
@@ 5.1 Task Quality and Embedding Evaluation Reliability
Our analysis reveals a striking pattern: models achieve their highest relative performance precisely where human experts show the least agreement. This confirms that on low-quality datasets, current metrics do not measure semantic understanding but rather the model’s ability to reproduce consistent annotation artifacts.
This finding reframes the role of human evaluation in benchmark design. Rather than serving merely as a performance benchmark, human consensus establishes a validity threshold for evaluation tasks. When models significantly exceed this bound on low-agreement tasks, it signals that the benchmark itself has lost its descriptive power: the task may be measuring annotation artifacts rather than the capability it claims to assess. HUME provides the empirical mechanism to identify and deprecate these invalid evaluation targets, ensuring that leaderboards measure genuine capability rather than overfitting to noise.
Tasks with high human agreement, such as reranking and toxicity classification, provide reliable evaluation targets. Conversely, low-agreement tasks (e.g., ArXiv clustering, Emotion classification) suffer from ambiguous guidelines or subjective judgments. Cultural factors add another dimension to evaluation reliability. Humans retain substantial advantages in Arabic semantic similarity and multilingual sentiment analysis, revealing genuine model limitations in cross-cultural understanding.
These findings suggest reliable evaluation depends as much on task quality as model capability. Rather than treating high model performance as automatic progress, we recommend prioritizing high-agreement tasks for development, addressing cultural competence gaps, and critically examining whether apparent model superiority on ambiguous tasks reflects genuine capability or evaluation bias. Novel benchmarks should report human agreement alongside model scores: 85% accuracy on emotion classification (185% of human performance, $\kappa=0.39$) represents a fundamentally different achievement than 85% on reranking (97% of human performance, $\rho=0.75$).
5.2 LLM-Based Evaluation
Our LLM annotation experiments reveal important limitations for using LLMs as proxies for human judgment. While LLMs achieve competitive performance on classification tasks, a substantial gap emerges on reranking, where humans significantly outperform even the best LLMs. Notably, reranking tasks show strong inter-annotator agreement ($\rho=0.64$$0.85$), suggesting LLMs struggle with precisely the nuanced relevance judgments where human consensus is highest. This pattern contrasts with classification, where lower human agreement ($\kappa=0.24$$0.55$) coincides with near-parity between humans and LLMs.
This has important implications for benchmark development. The reranking gap suggests that LLMs may not reliably capture the semantic distinctions humans make on well-defined tasks, even as they approach human performance on more ambiguous ones. Using LLMs for large-scale annotation may therefore introduce systematic biases, particularly for tasks requiring fine-grained semantic judgments. The architectural mismatch between generative LLMs and discriminative evaluation tasks further limits their utility, as evidenced by our inability to evaluate clustering tasks.
While LLMs offer scalability advantages, these limitations suggest they should augment rather than replace human annotation, particularly for benchmark development where task quality directly impacts model development priorities. Future work should explore hybrid approaches that leverage LLM efficiency for initial annotation while reserving human judgment for high-agreement tasks and uncertain cases.
@@ 5.3 Implications for Model Development and Evaluation Practices
Address Cultural and Linguistic Competence Gaps: The substantial human advantages in non-English tasks reveal critical model limitations that scaling training data alone cannot address. Arabic semantic similarity shows the largest human advantage (67.5% vs. 40.9% best model), while multilingual sentiment demonstrates consistent human superiority in non-English languages (95.0% Arabic, 92.5% Russian). These gaps suggest that current models lack the cultural and contextual knowledge necessary for cross-lingual understanding, requiring architectural innovations or training approaches that go beyond simple data scaling to capture cultural nuances and contextual understanding.
@@
Replace Problematic Benchmark Datasets: Our analysis identifies specific datasets that compromise benchmark reliability and should be replaced in future MTEB iterations: emotion classification , ArXiv clustering, and STS22-Russian (systematic parsing artifacts). These tasks provide unreliable evaluation targets that may mislead model development efforts by rewarding pattern matching to flawed gold standards. Replacement datasets should demonstrate reasonable human agreement and clear task specifications, validated through human evaluation before inclusion in benchmark suites.
Report Dataset Quality Measures: Model performance should be interpreted in light of dataset quality indicators to provide proper context for evaluation results. We propose that benchmark leaderboards report human agreement metrics alongside model scores. A model achieving 85% accuracy on emotion classification (185% of human performance, $\kappa=0.39$) represents a fundamentally different achievement than 85% on reranking (97% of human performance, $\rho=0.75$). High model performance on low-agreement tasks should be viewed skeptically as potential artifacts of flawed evaluation targets rather than genuine capability improvements. For tasks where human agreement falls below established thresholds ($\kappa<0.4$ or $\rho<0.6$), we recommend either improving task specifications or removing the dataset from benchmark suites entirely. However, it is important to recognize that some degree of human disagreement reflects natural variability in judgment rather than dataset flaws. Future benchmarks could incorporate evaluation frameworks that preserve and leverage this variability rather than collapsing it to single gold labels (plank-2022-problem; basile2020we).
5.4 Limitations
Our study has several limitations. First, our prioritization of breadth over depth—covering 16 diverse tasks—resulted in smaller sample sizes per task (20–50 instances). While we provide significance analyses to validate our statistical conclusions, larger samples would better capture the full complexity of human performance variation and provide more robust estimates of human judgment distributions.
Second, our multi-task design constrained task-specific training. Annotators were average or above-average raters without specialized training; experts would likely perform better, particularly on technical tasks. This was compounded by sparse annotation guidelines in original datasets, making alignment with original procedures difficult—though this reflects realistic annotation scenarios where perfect replication is often infeasible.
Third, while three annotators participated overall, only two annotations were collected for most tasks, limiting our ability to fully characterize agreement patterns. Additionally, while we ensured cultural and linguistic diversity among annotators, they were all male and aged 20–35, which does not fully represent human judgment distributions across broader demographic groups.
Fourth, While our study quantifies where models diverge from human performance, it does not fully explain why these gaps arise. Identifying the underlying factors - such as gaps in training data coverage, domain or cultural biases, and linguistic variability, particularly in low-resource settings - remains a critical direction for future research. However, detailed information about model training corpora is often unavailable, limiting such analysis.
Finally, while our study evaluates human performance across diverse tasks, we did not systematically investigate how task design features—such as specification clarity versus meaningful challenge—affect human agreement and model discrimination. The field needs rigorous research on these design principles to avoid both ambiguity, which depresses human agreement, and oversimplification, which diminishes discriminative power.
@@ 6 Conclusion
We introduce HUME, a comprehensive human evaluation framework for MTEB, addressing a critical gap in understanding empirical performance bounds for embedding models. By measuring human performance across 16 datasets spanning reranking, classification, clustering, and STS, we establish statistically robust baselines that reframe how model achievements should be interpreted.
Our findings show that human performance varies substantially by task categories. Tasks with high agreement provide reliable benchmarks, while low-agreement tasks often reveals design issues in the task formulation.
Finally, our benchmarking of nine LLM-as-annotator systems demonstrates that while they offer scalability, they cannot yet replace human judgment entirely. The best LLM (GPT-4.1-mini, 76.1%) falls short of human performance (81.2%), particularly on reranking. This suggests that future benchmarks should leverage LLMs to augment, but not replace, human evaluation.
6.1 Ethical considerations
Annotators were co-authors who consented to the study. There were no external crowd workers involved in any part of the annotation process.
Appendix A Dataset Specifications
Table 3 details the 16 datasets selected for this study, including their domains, descriptions, and the primary metrics used for evaluation.
Datasets Description Metrics
Reranking Core17, News21, Robust04 (weller2024followir) Information retrieval benchmarks (news, documents) MAP, MRR@10, nDCG@10
WikiMulti (enevoldsen2025mmteb) Wikipedia article reranking (eng, dan, nob)
Classification Emotion (saravia-etal-2018-carer) Emotion classification from social media text Accuracy, F1, Weighted F1
Tweet Senti (barbieri-etal-2022-xlm) Sentiment analysis of tweets
Toxicity (jigsaw-unintended-bias-in-toxicity-classification) Toxic content detection
Multilingual Sentiment (mollanorozy-etal-2023-cross) Sentiment classification (ara, eng, nob, rus)
Clustering WikiCities (wikidump) Entity clustering from Wikipedia V-Measure, ARI, AMI
ArXiv (arxiv_org_submitters_2024) Academic paper topic clustering (derived labels)
Reddit (geigle:2021:arxiv) Forum discussion topic clustering
SIB200 (adelani2023sib) Multilingual sentence clustering (ara, dan, eng, rus)
STS STSBenchmark (huggingface:dataset:stsb_multi_mt) General semantic similarity benchmark Spearman, Pearson
SICK-R (marelli-etal-2014-sick) Semantic relatedness and entailment
STS12 (Agirre2012sts) Shared task semantic similarity
STS22 (chen-etal-2022-semeval) Multilingual semantic similarity (ara, eng, rus)
Table 3: Complete list of 16 datasets and evaluation metrics used for human annotation. For the metrics we use MAP (Manning_Raghavan_Schütze_2008), MRR (Manning_Raghavan_Schütze_2008), nDCG (10.1145/582415.582418), Accuracy/F1 (SOKOLOVA2009427), V-Measure (rosenberg-hirschberg-2007-v), ARI (hubert1985comparing), AMI (article), Spearman/Pearson (10.1093/ije/dyq191) (Pearson1895), following the MTEB implementations.
Appendix B Detailed Results by Task Category
@@
Appendix C Task Examples
@@
Appendix D Task Quality Analysis
D.1 Dataset Quality Issues
@@ D.1.1 STS22-Russian
D.1.2 Multilingual sentiment classification-Russian
@@
  • Similar to the issues described in § D.1.1, the dataset contains parsing artifacts from web pages.
D.1.3 Emotion Classification
@@
D.1.4 ArXiv Clustering Challenges
@@ D.2 Impact on Evaluation
Appendix E Statistical Robustness Analysis
E.1 Confidence Interval Methodology
Given sample size constraints (N=2050N=20-50), we computed 95% Confidence Intervals (CIs) for human performance using metric-specific analytical methods rather than generic approximations:
  • Classification (Accuracy): Wilson Score Interval (Wilson:1927xyh), which is robust for binomial proportions with small sample sizes and avoids the “zero-width” errors of normal approximations.
  • STS (Correlation): Fisher $z$-transformation (10.1093/biomet/10.4.507) to compute CIs for Spearman correlations, ensuring valid bounds within [1,1][-1,1].
  • Clustering & Reranking: Empirical range between annotators as a conservative bound given $N_{\text{annotators}}=2$.
Statistical significance () is determined by a non-parametric overlap test: a model is considered significantly different if its score falls outside the human 95% CI (corresponding to $p<0.05$).
E.2 Detailed Results with Confidence Intervals
Table 8 presents human performance with 95% confidence intervals for all 26 task-language pairs. Models perform outside the human confidence interval in 14 of 26 tasks, indicating statistically significant differences. Crucially, this separation often occurs in tasks with low to moderate inter-annotator agreement. For example, in EmotionClassification, the best model (75.4) significantly exceeds human performance (45.8, CI:[32.6, 59.7]), but the low agreement ($\kappa=0.39$) suggests this “superhuman” performance may reflect fitting to annotation artifacts rather than genuine semantic superiority. Conversely, in MultilingualSentiment (Arabic) and STS22 (Arabic), humans significantly outperform models, highlighting genuine cultural gaps that current models fail to bridge.
Task Lang N K Tot Human (95% CI) IAA Best Model
Classification
EmotionClassification eng 48 2 96 45.8 [32.6, 59.7] $\kappa=0.39$ 75.4
MultilingualSentiment ara 40 1 40 95.0 [83.5, 98.6] N/A 77.5
MultilingualSentiment eng 40 2 80 77.5 [62.5, 87.7] $\kappa=0.24$ 95.5
MultilingualSentiment nob 40 1 40 85.0 [70.9, 92.9] N/A 75.0
MultilingualSentiment rus 40 1 40 92.5 [80.1, 97.4] N/A 81.3
ToxicConversations eng 45 2 90 73.3 [59.0, 84.0] $\kappa=0.55$ 86.7
TweetSentimentExtraction eng 45 2 90 84.4 [71.2, 92.3] $\kappa=0.41$ 90.9
Clustering
ArxivClusteringP2P eng 30 2 60 49.2 [35.3, 63.2] ARI=-0.00 84.6
RedditClusteringP2P eng 30 2 60 68.8 [63.2, 74.4] ARI=0.42 100.0
SIB200ClusteringS2S ara 30 1 30 76.0 [58.4, 87.8] N/A 78.8
SIB200ClusteringS2S dan 30 1 30 62.7 [44.9, 77.6] N/A 76.0
SIB200ClusteringS2S eng 30 2 60 54.0 [41.8, 66.3] ARI=0.15 83.3
SIB200ClusteringS2S rus 30 1 30 68.1 [50.2, 81.9] N/A 77.7
WikiCitiesClustering eng 30 2 60 97.6 [95.2, 100.0] ARI=0.91 100.0
Reranking
Core17Instruction eng 20 2 40 85.2 [83.6, 86.8] $\rho=0.80$ 98.8
News21Instruction eng 31 2 62 92.7 [91.3, 94.1] $\rho=0.85$ 100.0
Robust04Instruction eng 49 2 98 88.5 [82.2, 94.8] $\rho=0.75$ 98.8
WikipediaMultilingual dan 30 1 30 91.4 [76.2, 97.3] N/A 95.0
WikipediaMultilingual eng 30 2 60 82.4 [75.6, 89.1] $\rho=0.64$ 90.6
WikipediaMultilingual nob 30 1 30 89.8 [74.1, 96.4] N/A 92.3
STS
SICK-R eng 40 2 80 82.7 [69.4, 90.5] $\rho=0.63$ 94.1
STS12 eng 50 2 100 91.2 [84.9, 94.9] $\rho=0.77$ 92.0
STS22 ara 30 1 30 67.6 [41.7, 83.3] N/A 40.9
STS22 eng 30 2 60 78.4 [59.0, 89.2] $\rho=0.75$ 82.9
STS22 rus 30 1 30 58.7 [28.8, 78.2] N/A 69.5
STSBenchmark eng 50 2 100 80.4 [67.7, 88.4] $\rho=0.58$ 90.9
[…]
Appendix F Inter-Annotator Agreement Analysis
F.1 Agreement Metrics by Task Category
@@ F.1.4 STS Tasks
F.2 Agreement Patterns and Task Reliability
F.2.1 High-Agreement Tasks (Reliable Benchmarks)
@@
F.2.2 Low-Agreement Tasks (Problematic Benchmarks)